HELP!! I just need help with those two... Can you please help me?!?!?
Solve each system by elimination.
1)2x-y+z=-2 ~~~~ 2)x-2y+3z=12
x+3y-z=10 ~~~~ 2x-y-2z=5
x +2z=-8 ~~~~ 2x+2y-z=4

Answer by stanbon(75887)

Solve each system by elimination.
1)
2x-y+z=-2
x+3y-z=10
x+0y+2z=-8
----------------
Let the 3rd equation be the 1st.
x+0y+2z=-8
2x-y+z=-2
x+3y-z=10
-------------
2 times the 1st subtracted from the 2nd
Subtract the 1st from the 3rd
x+0y+2z=-8
0x-y-3z=14
0x+3y-3z=18
--------------------
Add 3 times the 2nd to the 3rd
x+0y+2z=-8
0x-y-3z=14
0x+0y-12z=60
--
Solve the 3rd to get: z = -5
Substitute into the 2nd to get y = 1
Substitute those values into the 1st to get x = 2
=======================================================

2)
x-2y+3z=12
-2x-y-2z=5
2x+2y-z=4
----------------
2 times the 1st added to the 2nd
2 times the 1st subtracted from the 3rd
x-2y+3z=12
0x-5y+4z=29
0x+6y-7z=-20
------------------------
Add the 2nd to the 3rd
x-2y+3z=12
0x-5y+4z=29
0x+y-3z=9
---------------------
Add 5 times the 3rd to the 2nd:
x-2y+3z=12
0x+0y-11z=74
0x+y-3z=9
---------
Solve the 2nd to get: z = -74/11
Substitute into the 3rd to get y = -123/11
Substitute into the 1st to get x = 108/11
==========================================
Cheers,
Stan H.